OpenStudy (anonymous):

use the quadratic formula to find the zeros y=2x^2+8x+1

OpenStudy (mathmate):

Quadratic formula to be used is: \( x=\frac{-b\pm\sqrt{n^2-4ac} }{2a} \) where a,b,c refer to an equation \( ax^2+bx+c=0 \) Here a=2,b=8,c=1. Can you find x?

OpenStudy (anonymous):

can you help more

OpenStudy (mathmate):

Substitute a=2, b=8, c=1 into the quadratic formula to get: \( x=\frac{-8\pm \sqrt{8^2-4(2)(1)}}{2(2)} \) Can you then evaluate that with a calculator or by hand?

OpenStudy (anonymous):

is there more to this problem, not getting an answer that i have to choose from

OpenStudy (mathmate):

What did you come up with, and what are the choices?